Circular No. 6270

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE The Supernova Cosmology Project [S. Perlmutter, S. Deustua, G.

Goldhaber, D. Groom, I. Hook, A. Kim, M. Kim, J. Lee, J. Melbourne, C. Pennypacker, and I. Small, Lawrence Berkeley Lab. and the Center for Particle Astrophysics; A. Goobar, Univ. of Stockholm; R. Pain, CNRS, Paris; R. Ellis and R. McMahon, Inst. of Astronomy, Cambridge; and B. Boyle, P. Bunclark, D. Carter, and M. Irwin, Royal Greenwich Obs.; with A. V. Filippenko and A. Barth (Univ. of California, Berkeley) at the Keck telescope; W. Couch (Univ. of N.S.W.) and M. Dopita and J. Mould (Mt. Stromlo and Siding Spring Obs.) at the Siding Spring 2.3-m telescope; H. Newberg (Fermi National Accelerator Lab.) and D. York (Univ. of Chicago) at the ARC telescope] report eleven supernovae found with the Cerro Tololo (CTIO) 4-m telescope in their 1995 High Redshift Supernova Search:

SN  1995 UT  α_{2000}  δ_{2000}  R  Offset
1995aq Nov. 19 0^{h}29^{m}04.^{s}26 + 7^{o}51′20′′.0 22.4 0′′.6 W, 1′′.4 S
1995ar Nov. 19 1 01 20.41 + 4 18 33.8 23.1 2′′.9 W, 0′′.5 S
1995as Nov. 19 1 01 35.30 + 4 26 14.8 23.3 0′′.7 W, 0′′.7 N
1995at Nov. 20 1 04 50.94 + 4 33 53.0 22.7 0′′.3 W, 0′′.4 S
1995au Oct. 29 1 18 32.60 + 7 54 03.5 20.7 1′′.4 E, 3′′.3 N
1995av Nov. 20 2 01 36.75 + 3 38 55.2 20.1 0′′.2 W, 0′′.0 N
1995aw Nov. 19 2 24 55.54 + 0 53 07.5 22.5 0′′.2 W, 0′′.2 S
1995ax Nov. 19 2 26 25.80 + 0 48 44.2 22.6 0′′.3 W, 0′′.2 S
1995ay Nov. 20 3 01 07.52 + 0 21 19.4 22.7 0′′.9 W, 1′′.4 S
1995az Nov. 20 4 40 33.59 − 5 30 03.6 24.0 1′′.6 W, 1′′.7 N
1995ba Nov. 20 8 19 06.46 + 7 43 21.2 22.6 0′′.1 E, 0′′.2 N

The spectra (Keck, Nov. 26–28) are consistent with type-I supernovae (except SN 1995av, a probable type II) at the redshift of the host galaxy: z = 0.45, 0.46, 0.49 (preliminary type-I identification), 0.65, 0.16, 0.30, 0.4 (supernova redshift only), 0.61, 0.48, 0.45, 0.39. Photometry obtained on Nov. 21–23 at CTIO (A. Walker) and Nov. 23–27 at WIYN (D. Harmer, D. Willmarth) indicates that SNe 1995ar, 1995at, 1995av, 1995aw, 1995ay, and 1995az are now before or at maximum, while the others are slightly past maximum. The previous observations not showing the supernovae (to limiting mag ∼ 24) were on Oct. 29–30 at the CTIO 4-m (except SN 1995au, on 1994 Sept. 29 at the Kitt Peak 4-m telescope). Continuing R, I, and B photometry is important. Contact saul@LBL.gov for finding charts.

1995 December 6 Daniel W. E. Green